OpenStudy (anonymous):

If h(x) = 3x - 4, find each of the following: a) h(2) b) h(a + 2)

OpenStudy (anonymous):

Put value 2 for x and solve it Same idea x=a+2 and solve it

OpenStudy (anonymous):

.... i still dont get it

OpenStudy (anonymous):

you see where h(x) is? basically for part a and b, all your doing is substituting part a in where the h(x) is and solve for it, same thing with part b. it might look like this 2 = 3x-4 (a+2)=3x-4 For part all just add the 4 to the to and then divide the x.

OpenStudy (anonymous):

f(2)=3(2)-4=6-4=2 f(a+2)=3(a+2)-4 = 3a+6-4 =3a+2